A wound is healing in such a way that t days since Monday, the area of the wound has been decreasing at a rate of -3(t+2)^-2 sq cm per day. If on Tuesday, the area of the wound was 2 sq cm, what was the area of the wound on Monday?
1
Expert's answer
2020-11-02T14:40:42-0500

"f(t) =-\frac{3}{(t+2)^2}"

"f(tuesday) = f(1) = -\frac{3}{(1+2)^2}=-\frac{1}{3}"

On Tuesday the wound was 2 sq sm and decreased at 1/3 sq sm since Monday.

Therefore on Monday the wound was:

"2+\frac{1}{3}=2.33 sm^2"